# bibliography.bib

### v1

@@ Abstract
Patterns suggesting the existence of a categorification of the homomorphism $\varpi_{\textrm{alg}}$ play a prominent role in the study of categorified link invariants, see [gnr]. This was one of the starting points of our investigation. The objects $\wedge^{k}\mathcal{S}_{n}$ (more precisely, their projection to the category of unipotent character sheaves), are objects representing Khovanov-Rozansky homology in the setting of monodromic Hecke categories, see [hhh].
In the setting of matrix factorizations, categorification of $\varpi_{\textrm{br}}$ was constructed by Oblomkov and Rozansky in [oblomkovAFFINEBRAIDGROUP2019].
Most of the results and ideas of this paper are contained in the author’s thesis [tolmachovFunctorAffineFinite2018]. The main difference is the replacement of the monoidal derived category of character sheaves with the abelian category, and technical complications arising from this, see the discussion in Section 3.2. In the note [bezrukavnikovTolmachovExt], joint with Bezrukavnikov, we consider a simpler analogue of some of the results presented here in the setting of the derived category of sheaves on the Lie algebra $\mathfrak{gl}_{n}$.
A categorification of $\varpi_{\textrm{alg}}$ whose source is an additive diagrammatic affine Hecke category is constructed in [MACKAAY2024109401]. See also [eliasGaitsgoryCentralSheaves2018] for the discussion of the diagrammatic approach to the triangulated categorification $\varpi_{\textrm{alg}}$.
A different approach to constructing a functor from the affine Hecke category is developed in [taoAffineHeckeCategory2021], which presents the monoidal $\infty$-version of the Hecke category as a monoidal colimit of the finite type Hecke categories associated to parahoric subgroups. The description of $\varpi_{\textrm{alg}}$ in terms of braid groups provides a natural candidate for the restriction of its categorification to these finite type Hecke categories. See [tao_talk] for the discussion of the compatabilities between these restrictions required to extend them to the affine Hecke category.
@@
Author is supported by the EPSRC programme grant EP/R034826/1.
@@
Let $G$ be an arbitrary reductive group. Let $\mathcal{H}^{\prime}(G)$ be the full triangulated subcategory of $D^{b}(U\backslash G/U)$ generated by the image of the functor $\pi^{*}:D^{b}(U\backslash G/B)\to D^{b}(U\backslash G/U),$ where $\pi$ is the map induced by the standard projection $G/U\to G/B$. Complexes in $\mathcal{H}^{\prime}(G)$ are characterized among all constructible sheaves on $U\backslash G/U$ by the property that their restriction to the fibers of the $T$-torsor $G/U\to G/B$ have smooth cohomology with unipotent monodromy.
In [bezrukavnikovKoszulDualityKacMoody2014], a completion $\mathcal{H}(G)\subset\operatorname{pro}\mathcal{H}^{\prime}(G)$ was defined. This is a triangulated monoidal category, equipped with a convolution operation $\star$, coming with three collections of objects $\hat{\Delta}_{w},\hat{\nabla}_{w},\hat{\mathcal{T}}_{w},$ indexed by the elements $w$ in the Weyl group $W$. These objects are referred to as pro-unipotent standard, costandard and tilting objects, respectively.
@@
    $\hat{\Delta}_{v}\star\hat{\Delta}_{w}\simeq\hat{\Delta}_{vw},\hat{\nabla}_{v}\star\hat{\nabla}_{w}\simeq\hat{\nabla}_{vw}$ if $l(vw)=l(v)+l(w)$.
@@
    $\hat{\Delta}_{w}\star\hat{\mathcal{T}}_{w_{0}}\simeq\hat{\mathcal{T}}_{w_{0}}\star\hat{\Delta}_{w}\simeq\hat{\mathcal{T}}_{w_{0}},\hat{\nabla}_{w}\star\hat{\mathcal{T}}_{w_{0}}\simeq\hat{\mathcal{T}}_{w_{0}}\star\hat{\nabla}_{w}\simeq\hat{\mathcal{T}}_{w_{0}},$
@@
(a) and (b) are [bezrukavnikovKoszulDualityKacMoody2014][Lemma 4.3.3], (c) is [bezrukavnikovTopologicalApproachSoergel2018][Lemma 7.7], (d) is [bezrukavnikovTopologicalApproachSoergel2018][Proposition 7.10]. ∎
Let $\mathcal{H}^{\mathrm{perf}}(G)$ stand for the full triangulated subcategory of $\mathcal{H}(G)$ generated by the object $\hat{\mathcal{T}}_{w_{0}}$.
@@
By the results of [bezrukavnikovKoszulDualityKacMoody2014], we have an equivalence of categories $\mathcal{H}(G)\simeq\operatorname{Ho}^{b}(\hat{\mathcal{T}})$, the bounded homotopy category of the full additive subcategory $\hat{\mathcal{T}}$ of $\mathcal{H}(G)$ consisting of direct sums of tilting objects. $\mathcal{H}^{\mathrm{perf}}(G)$ is the subcategory of complexes whose terms are direct sums of copies of the object $\hat{\mathcal{T}}_{w_{0}}$.
2.2   Jucys-Murphy sheaves.
Let $G=\textrm{GL}_{n}$ and write $\mathcal{H}_{n}=\mathcal{H}(G),\mathcal{H}_{n}^{\mathrm{perf}}=\mathcal{H}^{\mathrm{perf}}_{n}(G)$. Let $B(W)$ stand for the Artin-Tits braid group attached to the Coxeter group $W$. By Proposition 2.1.1, $\mathcal{H}_{n}$ admits left and right $B(W)$-actions by convolution with complexes of sheaves $F_{\beta},\beta\in B(W)$, defined on generators of $B(W)$ as follows. Let $w$ also stand for the standard lift of $w\in W$ to $B(W)$. Then we define $F_{w}:=\hat{\Delta}_{w},F_{w}^{-1}:=\hat{\nabla}_{w^{-1}}$. For an arbitrary braid $\beta$, $F_{\beta}$ is defined as the convolution of complexes corresponding to the generators. Define the braids $j_{k},k=1,\dots,n$ by $j_{1}=e,j_{k+1}=s_{k}j_{k}s_{k}$. The braids $j_{k}$ are called Jucys-Murphy braids, and it is well known that they generate a commutative subgroup of $B(W)$. For a weight $\lambda$ we denote by $\lambda$ the corresponding object in $\operatorname{Rep}(\mathbf{T})$. The assignment
$\lambda=\sum a_{i}e_{i}\mapsto\mathfrak{J}_{\lambda}=F_{j_{1}}^{\star a_{1}}\star\dots\star F_{j_{n}}^{\star a_{n}}$
defines left and right actions of $\operatorname{Rep}(\mathbf{T})$ on $\mathcal{H}_{n}$ by left and right convolution with $\mathfrak{J}_{\lambda}$, respectively. By Proposition 2.1.1 (d) this action preserves the subcategory $\mathcal{H}_{n}^{\mathrm{perf}}$.
@@
We will use extensively some results of [gabberFaisceauxPerversLadiques1996]. Let $T$ be an algebraic torus of rank $r$ defined over $\mathbf{k}$. Let $\pi^{t}_{1}(T)$ stand for the tame fundamental group of $T$ and $\pi_{1}^{\ell}(T)$ for its maximal $-\ell$ quotient. Let
@@
By Proposition 3.2.2 of [gabberFaisceauxPerversLadiques1996] $\mathfrak{R}$ is a regular Noetherian ring. Let
@@
where $\chi$ runs through the characters $\chi:\pi^{t}_{1}(T)\to\overline{\mathbf{Q}}_{\ell}^{\times}$ of finite order prime to $\ell$. Let $\mathbf{1}$ stand for the trivial character. $\overline{\mathbf{Q}}_{\ell}$-points of $\mathcal{C}(T)$ ($\mathcal{C}(T)_{\ell}$) are identified with characters $\pi_{1}^{t}(T)\to\overline{\mathbf{Q}}_{\ell}^{\times}$ (with continuous characters $\pi_{1}^{\ell}(T)\to\overline{\mathbf{Q}}_{\ell}^{\times}$). The group $\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$ acts on the scheme $\mathcal{C}(T)$ by automorphisms. Denote by $m_{\lambda}$ the automorphism corresponding to $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$.
@@
Recall that in [gabberFaisceauxPerversLadiques1996] the Mellin transform functor
@@
Let $\mathfrak{m}_{0}$ be the ideal of $\mathfrak{R}$ generated by $t_{1},\dots,t_{r}$. We regard $\mathfrak{R}/\mathfrak{m}_{0}^{n}$ as coherent sheaf on $\mathcal{C}(T)$ set-theoretically supported on a point inside the connected component $\{\mathbf{1}\}\times\mathcal{C}(T)_{\ell}.$ We record the following simple
@@
Let $\mathbf{N}$ for the category whose objects are non-negative integers, with $\operatorname{Hom}(i,j)$ having a single arrow if and only if $i\geq j$. Let $\mathcal{A}$ be an additive category and let $\mathcal{F}:\mathbf{N}\to\mathcal{A}$ be an inverse system in $\mathcal{A}$. Such system is called essentially zero (short for “essentially constant equal to zero”) if, for any $m\in\mathbb{N}$, there is $n>m$ such that $\mathcal{F}(n\rightarrow m)=0$.
@@
    System ${}^{p}\mathcal{H}^{i}(\mathcal{F}\star\mathcal{E}_{n}^{\lambda}[r])$ is essentially zero for all $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$ and all $i<0$.
@@
Note that since $\mathfrak{R}$ is regular Noetherian and the automorphism group acts transitively on the set of $\overline{\mathbf{Q}}_{\ell}$-points, $\mathcal{C}(T)_{\ell}$ has finite global dimension (see [stacks-project, Tag 00OE]), and $\mathscr{M}(\mathcal{F})|_{\mathbf{1}\times\mathcal{C}(T)_{\ell}}$ has a finite projective resolution. It follows that $\mathcal{H}^{i}(\mathscr{M}(\mathcal{F})\otimes^{\operatorname{L}}(\mathfrak{R}/\mathfrak{m}_{0}^{n}))$, considered as $\mathfrak{R}$-modules, are finite-dimensional over $\overline{\mathbf{Q}}_{\ell}$. From the short exact sequences
@@
Again, since the system in question is of finite-dimensional vector spaces, this implies that it is essentially zero. By property M5, we get that ${}^{p}\mathcal{H}^{i}(\mathcal{F}\star\mathcal{E}_{n}^{\lambda}[r])$ is essentially zero.
@@
Let $S\subset\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$ be a finite subset. We say that $\mathcal{F}\in D^{b}(T)$ is $S$-monodromic, if it is a direct sum of $\lambda$-monodromic complexes with $\lambda\in S$ (note that there are no morphisms between monodromic complexes with different monodromies).
@@
    System ${}^{p}\mathcal{H}^{i}(\mathcal{F}\star\mathcal{E}_{n}^{\lambda}[r])$ is essentially zero for all $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell}),\lambda\notin S$, and all $i$ if and only if ${}^{p}\mathcal{H}^{i}(\mathcal{F})$ are shifted $S$-monodromic local systems on $T$.
@@
    If for $\mathcal{F}\in P(T)$ we have ${}^{p}\mathcal{H}^{0}(\mathcal{F}\star\mathcal{L}_{\lambda}[r])=0$ for all $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell}),\lambda\notin S$, then $\mathcal{F}$ is a shifted $S$-monodromic local systems on $T$.
@@
Here we recall the necessary constructions from [boyarchenkoCharacterSheavesUnipotent2013] and some notations from the Introduction.
Let $G$ be a connected reductive group defined over $\mathbf{k}$. The category $D^{b}_{G}(G),$ denoted, where $G$ acts by conjugation, is bradied monoidal. We denote its braided structure by
@@
We also consider the “underived” convolution $\star^{0}$ on $P_{G}(G)$, devined as
@@
This equips $P_{G}(G)$ with the structure of braided monoidal category.
For any $M\in D^{b}_{G}(G)$ we have an isomorphism $\theta_{M}:M\to M$ called a twist, see loc.cit. It satisfies
$\theta_{A\star B}=\beta_{A,B}\circ\beta_{B,A}\circ(\theta_{A}\star\theta_{B})$ (3)
On the stalk of $M$ at a point $g\in G$ it acts by the conjugation by $g\in Z_{G}(g)$ — centralizer of $g$ in $G$.
@@
Since any $g\in G$ lies in $Z_{G}(g)^{0}$, automorphism $\theta$ acts as identity on stalks of any equivariant complex $\mathcal{F}\in D^{b}_{G}(G)$. This implies that $\theta$ is identity on any perverse sheaf, which, by (3), gives the result. ∎
@@
Note that in [tolmachovFunctorAffineFinite2018] the braided structure on $D^{b}_{\textrm{GL}_{n}}(\textrm{GL}_{n})$ was mistakenly treated as symmetric. Current, more technically involved, paper, thus, closes the gap in [tolmachovFunctorAffineFinite2018]. We expect, however, that the construction in [tolmachovFunctorAffineFinite2018] can be fixed also by modifying the braided structure on the derived category $D^{b}_{\textrm{GL}_{n}}(\textrm{GL}_{n})$, see Subsection 3.2.1 below.
@@
Let $(\mathcal{C},\star,\beta)$ be any braided monoidal category that is $k$-linear over a field $k$ of characteristic 0. Under certain conditions on the braided monoidal category $\mathcal{C}$, the techniques of [drinfeldQuasitriangularQuasiHopfAlgebras1991] (see also [etingofQuasisymmetricUnipotentTensor2009]) can be used to modify the braided structure on $\mathcal{C}$ to make it symmetric. Namely, say that the braided structure on $\mathcal{C}$ is unipotent if, for any triple of objects $A,B,C\in\mathcal{C}$ the subalgebra in $\operatorname{End}_{\mathcal{C}}(A\star B\star C)$ generated by $((\beta_{B,A}\circ\beta_{A,B})\star\operatorname{id}_{C}-1)$ and $(\operatorname{id}_{A}\star(\beta_{C,B}\circ\beta_{B,C})-1)$ is nilpotent. We put forward the following conjecture.
@@
It is easy to see, see Section 3.2, that if any element $g\in G$ lies in the neutral connected component of its centralizer in $G$, we have that $(\theta_{A}-1)$ is a nilpotent element in $\operatorname{End}(A)$ for any $A\in D^{b}_{G}(G)$. Since all the centralizers of elements of $\textrm{GL}_{n}$ are connected, this property holds for $\textrm{GL}_{n}$. In this case, the braiding being unipotent is equivalent to the condition that for any triple of objects $A,B,C\in\mathcal{C}$ the subalgebra in $\operatorname{End}_{\mathcal{C}}(A\star B\star C)$ generated by $(\theta_{A\star B}\star\operatorname{id}_{C}-1)$ and $(\operatorname{id}_{A}\star\theta_{B\star C}-1)$ is nilpotent.
@@
We mostly follow [bezrukavnikovParabolicRestrictionPerverse2021] in our geometric setup.
Let $G$ be a connected reductive group defined over $\mathbf{k}$, let $P\subset G$ be its parabolic subgroup, $U\subset P$ its unipotent radical and $L$ its Levi subgroup.
@@
where $L$ acts via the right diagonal action.
$G$ acts on $\mathcal{Y}_{P}$ and via the left diagonal action.
We equip the category $D^{b}_{G}(\mathcal{Y}_{P})$ with monoidal structure $\star$ given by *-convolution: to define it, we identify $D^{b}_{G}(\mathcal{Y}_{P})$ with $D^{b}_{U^{2}}(G/_{\operatorname{Ad}}L)$, where the subscript $U^{2}$ stands to denote full subcategory of complexes with constant cohomology along $U^{2}$-orbits, and use the formula (2).
@@
${G\times G/P}$${G}$${\mathcal{Y}_{P}}$$\scriptstyle{p}$$\scriptstyle{q}$
where $p$ is the projection and $q$ is the map $q(g,xP)=(xU,gxU)$. Consider the adjoint action of $G$ on itself. The functor
$\mathfrak{hc}=q_{*}p^{!}:D^{b}(G/G)\to D^{b}(\mathcal{Y}_{P}/G)$
is called the Harish-Chandra transform. It is well-known to be monoidal with respect to the *-convolution.
It has a left adjoint functor
$\chi=p_{*}q^{!}.$
For parabolics $P\subset Q$ we have functors $\mathfrak{hc}_{P}^{Q},\chi_{P}^{Q}$, defined in [lusztigParabolicCharacterSheaves2004]. We recall the definition. Let
$\tilde{Z}_{P,Q}=\{(xU_{Q},yU_{Q},zU_{P}):x^{-1}z\in Q\},$
and let $Z_{P,Q}$ be its quotient by $L_{Q}\times L_{P}$ action on the right. We have maps
@@
Write
$\chi_{P}^{Q}=f_{P,Q*}g_{P,Q}^{!}:D^{b}(\mathcal{Y}_{P}/G)\to D^{b}(\mathcal{Y}_{Q}/G),$

### v2

@@ Abstract
Patterns suggesting the existence of a categorification of the homomorphism $\varpi_{\textrm{alg}}:H_{n}^{\mathrm{aff}}\to H_{\mathfrak{n}}$ play a prominent role in the study of categorified link invariants [GNR21]. This was one of the starting points of our investigation. The (“mixed” analogues of) objects $\wedge^{k}\mathcal{S}_{n}$ (more precisely, their projection to the category of unipotent character sheaves) are objects representing Khovanov–Rozansky homology in the setting of monodromic Hecke categories [BT22].
In the setting of matrix factorizations, the categorification of $\varpi_{\textrm{br}}$ was constructed by Oblomkov and Rozansky in [OR19].
Most of the results and ideas of this paper are contained in the author’s thesis [Tol18]. The main difference is the replacement of the monoidal derived category of character sheaves with the abelian category, and technical complications arising from this. This replacement is needed because the derived category $D^{b}_{\textrm{GL}_{n}}(\textrm{GL}_{n})$ is not symmetric braided, but was mistakenly treated as such in ibid. We expect, however, that the braiding on the derived category can be modified to make it symmetric and make the simpler argument of [Tol18] go through. See the discussion in Section 3.2.
In the note [BT24], joint with Bezrukavnikov, we consider a simpler analogue of some of the results presented here in the setting of the derived category of sheaves on the Lie algebra $\mathfrak{gl}_{n}$.
A categorification of $\varpi_{\textrm{alg}}$ whose source is an additive diagrammatic affine Hecke category is constructed in [MMV24]. See also [Eli18] for a discussion of the diagrammatic approach to the triangulated categorification $\varpi_{\textrm{alg}}$.
A different approach to constructing a functor from the affine Hecke category is developed in [TT21], which presents the monoidal $\infty$-version of the Hecke category as a monoidal colimit of the finite type Hecke categories associated with parahoric subgroups. The description of $\varpi_{\textrm{alg}}$ in terms of braid groups provides a natural candidate for the restriction of its categorification to these finite type Hecke categories. See [Tao21] for the discussion of the compatabilities between these restrictions required to extend them to the affine Hecke category.
@@
I would like to thank the anonymous referees for a careful reading of the manuscript and many suggestions on improving the presentation.
The author was supported by the EPSRC programme grant EP/R034826/1 and by the Deutsche Forschungsgemeinschaft SFB 1624 grant, Projektnummer 506632645.
@@
[…]
In [BY13], a completion $\mathcal{H}(G)\subset\operatorname{pro}\mathcal{H}^{\prime}(G)$ was defined. This is a triangulated monoidal category, equipped with a convolution operation $\star$, coming with three collections of objects $\hat{\Delta}_{w},\hat{\nabla}_{w},\hat{\mathcal{T}}_{w},$ indexed by the elements $w$ in the Weyl group $W$. These objects are referred to as free-monodromic standard, costandard and tilting objects, respectively.
@@
    $\hat{\Delta}_{v}\star\hat{\Delta}_{w}\simeq\hat{\Delta}_{vw}$ and $\hat{\nabla}_{v}\star\hat{\nabla}_{w}\simeq\hat{\nabla}_{vw}$ if $l(vw)=l(v)+l(w)$.
@@
    $\hat{\Delta}_{w}\star\hat{\mathcal{T}}_{w_{0}}\simeq\hat{\mathcal{T}}_{w_{0}}\star\hat{\Delta}_{w}\simeq\hat{\mathcal{T}}_{w_{0}}$ and $\hat{\nabla}_{w}\star\hat{\mathcal{T}}_{w_{0}}\simeq\hat{\mathcal{T}}_{w_{0}}\star\hat{\nabla}_{w}\simeq\hat{\mathcal{T}}_{w_{0}}.$
@@
(a) and (b) are [BY13, Lemma 4.3.3], (c) is [BR22, Lemma 7.7], (d) is [BR22, Proposition 7.10]. ∎
Note that the equivalence of [Bez16] assigns to the structure sheaf $\mathcal{O}_{\mathbf{St}_{\mathfrak{g}}}$ the free-monodromic tilting object $\hat{\mathcal{T}}_{w_{0}}$ on the finite Hecke subcategory. It can be shown that the perfect derived category of $\mathbf{St}_{\mathfrak{g}}$ is generated, as an idempotent-complete triangulated category, by the two sided monoidal ideal generated by $\mathcal{O}_{\mathbf{St}_{\mathfrak{g}}}$. This motivates the following definition. Let $\mathcal{H}^{\mathrm{perf}}(G)$ stand for the full triangulated subcategory of $\mathcal{H}(G)$ generated by the object $\hat{\mathcal{T}}_{w_{0}}$. It is easy to deduce from Proposition 2.1.1 that $\mathcal{H}^{\mathrm{perf}}(G)$ is a two-sided monoidal ideal in $\mathcal{H}(G)$.
@@
By the results of [BY13], we have an equivalence of categories $\mathcal{H}(G)\simeq\operatorname{Ho}^{b}(\operatorname{Tilt}^{\wedge}(G))$, where the category on the right-hand side is the bounded homotopy category of the full additive subcategory $\operatorname{Tilt}^{\wedge}(G)$ of $\mathcal{H}(G)$ consisting of free-monodromic tilting objects. Under this equivalence, $\mathcal{H}^{\mathrm{perf}}(G)$ corresponds to the subcategory of complexes whose terms are direct sums of copies of $\hat{\mathcal{T}}_{w_{0}}$.
2.2  Jucys–Murphy sheaves and the $\operatorname{Rep}(\mathbf{T})$-action.
[…]
Define the braids $j_{k},k=1,\dots,n$ by $j_{1}=e,j_{k+1}=\tilde{s}_{k}j_{k}\tilde{s}_{k}$. The braids $j_{k}$ are called Jucys–Murphy braids, and it is well known that they generate a commutative subgroup of $B(W)$. We now return to the notations of the introduction, where $\mathbf{G}=\mathbf{GL}_{n}$ and $\mathbf{T}$ is a maximal torus of $\mathbf{G}$. For a weight $\boldsymbol{\lambda}$ we denote by $\boldsymbol{\lambda}$ the corresponding object in $\operatorname{Rep}(\mathbf{T})$. The assignment
$\boldsymbol{\lambda}=\sum a_{i}\mathbf{e}_{i}\mapsto\Phi_{l}(j_{1}^{a_{1}}\dots j_{n}^{a_{n}}),$
defines the left action of $\operatorname{Rep}(\mathbf{T})$ on $\mathcal{H}_{n}$. It follows from the discussion above, that the action of $\boldsymbol{\lambda}$ is isomorphic to the left convolution with an object
$\mathfrak{J}_{\boldsymbol{\lambda}}=F_{j_{1}}^{\star a_{1}}\star\dots\star F_{j_{n}}^{\star a_{n}}.$
The right action is defined analogously.
By Proposition 2.1.1 (d) this action preserves the subcategory $\mathcal{H}_{n}^{\mathrm{perf}}$.
@@
We will use extensively some results of [GL96]. Let $T$ be an algebraic torus of rank $r$ defined over $\mathbf{k}$. Let $\pi^{t}_{1}(T)$ stand for the tame fundamental group of $T$ and $\pi_{1}^{\ell}(T)$ for its maximal pro-$\ell$ quotient. Let
@@
By Proposition 3.2.2 of [GL96] $\mathfrak{R}$ is a regular Noetherian ring. Let
@@
where $\chi$ runs through the characters $\chi:\pi^{t}_{1}(T)\to\overline{\mathbf{Q}}_{\ell}^{\times}$ of finite order prime to $\ell$. Let $\mathbf{1}$ stand for the trivial character of $\pi_{1}^{t}(T)$. $\overline{\mathbf{Q}}_{\ell}$-points of $\mathcal{C}(T)$, resp. $\mathcal{C}(T)_{\ell}$, are identified with characters $\pi_{1}^{t}(T)\to\overline{\mathbf{Q}}_{\ell}^{\times}$, resp. with continuous characters $\pi_{1}^{\ell}(T)\to\overline{\mathbf{Q}}_{\ell}^{\times}$. The group $\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$ acts on the scheme $\mathcal{C}(T)$ by automorphisms. Denote by $m_{\lambda}$ the automorphism corresponding to $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$.
@@
Recall that in [GL96] the Mellin transform functor
@@
Let $\mathfrak{m}_{0}$ be the ideal of $\mathfrak{R}$ generated by $t_{1},\dots,t_{r}$. We regard $\mathfrak{R}/\mathfrak{m}_{0}^{n}$ as a coherent sheaf on $\mathcal{C}(T)$ set-theoretically supported on a point inside the connected component $\{\mathbf{1}\}\times\mathcal{C}(T)_{\ell}.$ We record the following simple
@@
Let $\mathbf{N}$ be the category whose objects are non-negative integers, with $\operatorname{Hom}(i,j)$ having a single arrow if and only if $i\geq j$. Let $\mathcal{A}$ be an additive category and let $\mathcal{F}:\mathbf{N}\to\mathcal{A}$ be an inverse system in $\mathcal{A}$. Such a system is called essentially zero (short for “essentially constant equal to zero”) if, for any $m\in\mathbb{N}$, there is $n>m$ such that $\mathcal{F}(n\rightarrow m)=0$.
@@
    The system of $^{p}\mathcal{H}^{i}(\mathcal{F}\star\mathcal{E}_{n}^{\lambda}[r])$ is essentially zero for all $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$ and all $i<0$.
@@
Note that since $\mathfrak{R}$ is regular Noetherian and the automorphism group $\mathcal{C}(T)_{\ell}(\overline{\mathbf{Q}}_{\ell})$ acts transitively on the set of $\overline{\mathbf{Q}}_{\ell}$-points, $\mathcal{C}(T)_{\ell}$ has finite global dimension (see [Sta22, Tag 00OE]), and $\mathscr{M}(\mathcal{F})|_{\mathbf{1}\times\mathcal{C}(T)_{\ell}}$ has a finite projective resolution. It follows that the cohomology $\mathfrak{R}$-modules $\mathcal{H}^{i}(\mathscr{M}(\mathcal{F})\otimes^{\operatorname{L}}(\mathfrak{R}/\mathfrak{m}_{0}^{n}))$ are finite-dimensional over $\overline{\mathbf{Q}}_{\ell}$. From the short exact sequences
@@
Again, since the system in question is of finite-dimensional vector spaces, this implies that it is essentially zero. By property M5, we get that the system of $^{p}\mathcal{H}^{i}(\mathcal{F}\star\mathcal{E}_{n}^{\lambda}[r])$ is essentially zero.
@@
Let $S\subset\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$ be a finite subset. We say that $\mathcal{F}\in D^{b}(T)$ is $S$-monodromic, if it is a direct sum of $\lambda$-monodromic complexes with $\lambda\in S$ (note that there are no non-zero morphisms between monodromic complexes with different monodromies).
@@
    The system of $^{p}\mathcal{H}^{i}(\mathcal{F}\star\mathcal{E}_{n}^{\lambda}[r])$ is essentially zero for all $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$ with $\lambda\notin S$, and all $i$, if and only if $^{p}\mathcal{H}^{i}(\mathcal{F})$ are shifted $S$-monodromic local systems on $T$.
@@
    If for $\mathcal{F}\in P(T)$ we have $^{p}\mathcal{H}^{0}(\mathcal{F}\star\mathcal{L}_{\lambda}[r])=0$ for all $\lambda\in\mathcal{C}(T)(\overline{\mathbf{Q}}_{\ell})$ with $\lambda\notin S$, then $\mathcal{F}$ is a shifted $S$-monodromic local system on $T$.
@@
Here we recall the necessary constructions from [BD14, Appendix B] and some notations from the introduction.
Let $G$ be a connected reductive group defined over $\mathbf{k}$. The category $D^{b}_{G}(G),$ where $G$ acts on itself by conjugation, is braided monoidal. We denote its braided structure by
@@
We also consider the “underived” convolution $\star^{0}$ on $P_{G}(G)$, defined as
@@
This equips $P_{G}(G)$ with the structure of a braided monoidal category.
For any $M\in D^{b}_{G}(G)$ we have an isomorphism $\theta_{M}:M\to M$ called a twist (see ibid). It satisfies
$\theta_{A\star B}=\beta_{B,A}\circ\beta_{A,B}\circ(\theta_{A}\star\theta_{B})$ (3)
Let $Z_{G}(g)$ stand for the centralizer of a point $g\in G(\mathbf{k})$. Since $M$ is equivariant with respect to the adjoint action, the stalk $M_{g}$ of $M$ at $g$ has an action of $Z_{G}(g)$. It is known that the action of $\theta_{M}$ on $M_{g}$ can be identified with the action of $g\in Z_{G}(g)(\mathbf{k}).$
@@
Since any $g\in G$ lies in $Z_{G}(g)^{0}$, automorphism $\theta$ acts by the identity on stalks of any equivariant complex $\mathcal{F}\in D^{b}_{G}(G)$. This implies that $\theta$ is the identity on any perverse sheaf, which, by (3), gives the result. ∎
@@
Note that in [Tol18] the braided structure on $D^{b}_{\textrm{GL}_{n}}(\textrm{GL}_{n})$ was mistakenly treated as symmetric. The current, more technically involved, paper, thus, closes the gap in [Tol18]. We expect, however, that the construction in [Tol18] can be fixed also by modifying the braided structure on the derived category $D^{b}_{\textrm{GL}_{n}}(\textrm{GL}_{n})$. See Subsection 3.2.1 below.
@@
Let $(\mathcal{C},\star,\beta)$ be any braided monoidal category that is $k$-linear over a field $k$ of characteristic 0. Under certain conditions on the braided monoidal category $\mathcal{C}$, the techniques of [Dri91] (see also [EG08]) can be used to modify the braided structure on $\mathcal{C}$ to make it symmetric. For a triple of objects $A,B,C\in\mathcal{C}$ consider the following endomorphisms of $A\star B\star C$:
$\nu_{A,B}$ $=(\beta_{B,A}\circ\beta_{A,B})\star\operatorname{id}_{C}-\operatorname{id}_{A\star B\star C},$
$\nu_{B,C}$ $=\operatorname{id}_{A}\star\,(\beta_{C,B}\circ\beta_{B,C})-\operatorname{id}_{A\star B\star C}.$
We say that the braided structure on $\mathcal{C}$ is unipotent if the subalgebra in $\operatorname{End}_{\mathcal{C}}(A\star B\star C)$ generated by $\nu_{A,B}$ and $\nu_{B,C}$ is nilpotent for any triple $A,B,C$. We put forward the following conjecture.
@@
It is easy to see (Section 3.2), that if any element $g\in G$ lies in the neutral connected component of its centralizer in $G$, we have that $\theta_{A}-\operatorname{id}_{A}$ is a nilpotent element in $\operatorname{End}(A)$ for any $A\in D^{b}_{G}(G)$. Since all the centralizers of elements of $\textrm{GL}_{n}$ are connected, this property holds for $\textrm{GL}_{n}$. In this case, the braiding being unipotent is equivalent to the condition that for any triple of objects $A,B,C\in\mathcal{C}$ the subalgebra in $\operatorname{End}_{\mathcal{C}}(A\star B\star C)$ generated by $\theta_{A\star B}\star\operatorname{id}_{C}-\operatorname{id}_{A\star B\star C}$ and $\operatorname{id}_{A}\star\,\theta_{B\star C}-\operatorname{id}_{A\star B\star C}$ is nilpotent.
@@
We mostly follow [BY21] in our geometric setup.
Let $G$ be a connected reductive group defined over $\mathbf{k}$, let $P\subset G$ be a parabolic subgroup, $U\subset P$ its unipotent radical and $L$ its Levi subgroup.
@@
where $L$ acts via the right diagonal action. Note that $G$ acts on $\mathcal{Y}_{P}$ via the left diagonal action.
We equip the category $D^{b}_{G}(\mathcal{Y}_{P})$ with the monoidal structure $\star$ given by *-convolution: to define it, we identify $D^{b}_{G}(\mathcal{Y}_{P})$ with $D^{b}_{U^{2}}(G/_{\operatorname{Ad}}L)$, where the subscript $U^{2}$ stands for the full subcategory of complexes with constant cohomology along $U^{2}$-orbits, and use the formula (2).
@@
[…]
where $p$ is the projection and $q$ is the map $q(g,xP)=(xU,gxU)L$. The functor
$\mathfrak{hc}=q_{*}p^{!}:D^{b}_{G}(G)\to D^{b}_{G}(\mathcal{Y}_{P}),$
where the action of $G$ on itself is the adjoint action, is called the Harish-Chandra transform. It is well-known to be monoidal with respect to the *-convolution. It has a left adjoint functor
$\chi=p_{!}q^{*}\simeq p_{*}q^{!}[-2\dim U].$
More generally, for parabolics $P\subset Q$, we have functors $\mathfrak{hc}_{P}^{Q},\chi_{P}^{Q}$, defined in [Lus04]. We recall the definition. Let
$\tilde{Z}_{P,Q}=\{(xU_{Q},yU_{Q},zU_{P})\in G/U_{Q}\times G/U_{Q}\times G/U_{P}:x^{-1}z\in Q\},$
and let $Z_{P,Q}=\tilde{Z}_{P,Q}/(L_{P}\times L_{Q})$, where $L_{P}$ acts on the $G/U_{P}$-factor on the right, and $L_{Q}$ acts diagonally on the right on the $G/U_{Q}\times G/U_{Q}$-factor. We have maps
@@
Note that $g_{P,Q}$ is well-defined because $x^{-1}z\in Q$, while $Q$ normalizes $U_{Q}$, and $U_{Q}\subset U_{P}$, since $P\subset Q$. Write
$\chi_{P}^{Q}$ $=f_{P,Q*}g_{P,Q}^{!}[-2\dim Q/P]:D^{b}(\mathcal{Y}_{P}/G)\to D^{b}(\mathcal{Y}_{Q}/G),$